### v3

# Adsorption of volatiles on dust grains in protoplanetary disks

Abstract
The adsorption of volatile molecules onto dust grain surfaces fundamentally influences dust-related processes, including condensation of gas-phase molecules, dust coagulation, and planet formation in protoplanetary disks. Using advanced ab-initio density functional theory with r2SCAN+rVV10 van der Waals functionals, we calculate adsorption energies of H2, H2O, and CO on carbonaceous (graphene, amorphous carbon) and silicate (MgSiO3) surfaces. Results reveal fundamentally different adsorption mechanisms: weak physisorption on carbonaceous surfaces ($|\Delta\epsilon_{\rm ad}|\sim 0.1-0.2~{\rm eV}$) versus strong chemisorption on silicates ($|\Delta\epsilon_{\rm ad}|\sim 0.5-1.5~{\rm eV}$) via coordination bonds. Kinetic Monte Carlo simulations incorporating these energies demonstrate divergent surface evolution: carbonaceous grains exhibit distinct condensation radius compared to silicates, while the cocrystal of H2O and CO significantly increases the desorption temperature of CO. The actual radii of gas-phase molecule depletion could thus be a comprehensive result of temperatures, chemical compositions, and even evolution tracks. Meanwhile, silicates maintain chemisorbed molecular coatings throughout most disk regions. Such dichotomy in surface coverage could also provide a natural mechanism for carbon depletion in inner planetary systems.
Protoplanetary disks(1300), Exoplanet formation (492), Interstellar dust (836), Dust physics (2229), Astrophysical dust processes (99)
I Introduction
The adsorption of volatile chemical species onto the surfaces of dust grains is of broad and critical interest to researchers studying the interstellar medium (ISM), star formation, and planet formation. For example, in protoplanetary disks (PPDs hereafter), dusts may play a central role in controlling the thermochemical conditions. Grains also hold the chemical equilibrium of formation versus dissociation inside disks, as their surfaces are the places where most $\mathrm{H_{2}}$ and $\mathrm{H_{2}O}$ are formed (e.g. Cazaux and Tielens, 2002, 2004; Ádámkovics et al., 2014). Moreover, dust surfaces can adsorb gas-phase charge carriers (ions and free electrons) rather efficiently, letting these carriers deposit their free charge and get neutralized as they get desorbed (e.g. Ilgner and Nelson 2006; see also Bai and Goodman 2009). This mechanism maintains a low but non-zero fraction of ionization in the pertinent zones in PPDs, enabling the non-ideal magnetohydrodynamic (MHD) effects to drive disk winds and accretion flows efficiently (Bai and Goodman, 2009; Bai, 2011, 2017; Wang et al., 2019). Volatile coatings on dust grain surfaces, for example water, regulate grain sintering and thereby shape the dust size distribution throughout PPD evolution (Okuzumi et al., 2016; Hasegawa et al., 2017). Such modulated distributions could influence their interactions with charged particles and lead to the formation of sub-structures in PPDs by alternating the magnetic diffusivities, the key parameters determining the interactions between the magnetic fields and the gas (e.g. Hu et al., 2019, 2021). More generally, the dynamic events related to dust growth, including coagulation, bouncing, breaking, and shattering of grains, are sensitively dependent on the sticking energy between surfaces (Chokshi et al., 1993; Dominik and Tielens, 1997). Such sticking energy should also be susceptible to the conditions on the grain surfaces. Discussions regarding these events are extensive under various circumstances, including theoretical analyses (e.g. Yan et al., 2004; Hirashita and Yan, 2009; Ormel et al., 2009; Hirashita and Voshchinnikov, 2014), and even experimental studies (Blum, 2000).
As the crucial characterization of adsorption processes, the energy of adsorption largely determines the status of volatile species and dust grains themselves in dust-rich astrophysical systems. It is widely assumed, based on observations and experiments, that the astrophysical dust grains can be broadly categorized into silicate grains and carbonaceous grains, whose mass fractions are in the same order of magnitude (see e.g. Jaeger et al. 1994; Dorschner et al. 1995; Jager et al. 1998; Fabian et al. 2000; Henning 2010; see also the data compilations in Draine 2011).
Over the past few decades, the adsorption condition of key species has been measured mainly by experimental studies. These experimental results should, nevertheless, be clarified and specified for their applicability and compatibility. For example, the reported adsorption energy of water on graphite is measured to be $E_{\mathrm{ad}}/k_{\mathrm{B}}\simeq 4800~{\rm K}$, almost identical to the adsorption on silicates (Sandford and Allamandola, 1990; Viti et al., 2004; Bolina et al., 2005; Ilgner and Nelson, 2006; Minissale et al., 2022). These values neveretheless seem at odds with the quantum chemistry studies concluding that graphite is only mildly hydrophilic by adsorbing molecules with the van der Waals (vdW hereafter) interactions (see also Ma et al., 2011; Hamada, 2012; Brandenburg et al., 2019), while the silicate surfaces could adsorb water mainly by covalent or coordination bonds and hydrogen bonds. Additionally, experimental studies have focused on the single-species adsorption conditions only, while the adsorbates in PPDs could be a mixture of multiple molecular species. The impact of the surface coverage conditions of grains deserve further and more thorough discussions with modern theoretical and computational methods.
In this paper, we will thoroughly study the problems related to the adsorption of molecules via computational chemistry. We will focus on the adsorption of water ($\mathrm{H_{2}O}$) and carbon monoxide ($\mathrm{CO}$), the most abundant molecules in the PPDs and molecular clouds (other than $\mathrm{H_{2}}$ and $\mathrm{He}$), onto the representative dust grain surface models, emphasizing the physical pictures of the results and compare them with experiments. This paper is structured as follows. §II describes the ab-initio calculation methods of adsorption on carbonaceous and silicate grains, assuming single and multiple molecule adsorption scenarios. Using these data, §III evaluates the surface coating conditions of both carbonaceous and silicate dust grains. §V summarizes the paper, and extends the scope of the results to the gas-phase molecule condensation and sublimation processes onto grain surfaces, the potential impact on planet formation mechanisms.
Figure 1: Top-down views of the model carbonaceous substrates studied in this article, including graphene (upper panel) and amorphous carbon (lower panel). The supercells used in the DFT calculations are illustrated by black solid lines in each panel, respectively.
II Ab-initio adsorption energy calculations
Figure 2: Left column: Adsorption energy as functions of molecular orientations and distances to the substrates for $\mathrm{CO}$ (upper row), $\mathrm{H_{2}O}$ (middle row), and $\mathrm{H_{2}}$ (lower row) molecules over graphene substrates. Right column: Illustrations of the molecular orientation notations. The depicted molecules correspond to the energy curve panels in the left column.
The condensation process of key molecules onto astrophysical grain surfaces is calculated in this work by examining adsorption conditions using advanced computational chemistry techniques. The ab-initio density functional theory (DFT) package VASP (Kresse and Hafner 1993; Kresse and Furthmüller 1996) is adopted, utilizing projector augmented wave (PAW) method calculations with PAW potentials for the atoms involved. To ensure accurate estimations of van der Waals (vdW) interactions, the combination of r2SCAN Sun et al. (2015); Furness et al. (2020) and rVV10 Sabatini et al. (2013) functionals is employed for electron exchange-correlation (XC) interactions. This combination, which involves generalized gradient approximations (GGA) and extensions to higher-order derivatives (meta-GGA), is confirmed by various tests and benchmarks (Peng et al., 2016) to handle vdW interactions under the DFT framework accurately in a wide range of situations, especially physisorptions. For instance, the single-molecule physisorption of water onto graphene, which is a highly relevant case to this work, is found to yield an adsorption energy $|\Delta\epsilon_{\mathrm{ad}}|\simeq 0.096~{\rm eV}$ when using r2SCAN+rVV10, a result that is remarkably close to fiducial quantum Monte Carlo (QMC) calculations (Ma et al. 2011; Brandenburg et al. 2019). Within the PAW paradigm, a set of plane-wave bases is utilized to expand the electronic wave functions, with a cutoff energy of $600~{\rm eV}$.
To measure the interaction energy of the systems, the molecule of interest is first positioned at the desired configuration to obtain the total energy with interaction, $\epsilon_{\mathrm{tot}}$. The molecule $X$ is then moved sufficiently far from the substrates to ensure that interactions between $X$ and the substrate (as well as neighboring adsorbates, in some cases) became negligible. The DFT calculation is repeated to obtain the sum of the energies of $X$ and the substrate, $\epsilon_{X}+\epsilon_{\mathrm{sub}}$. Based on the data collected, the adsorption energy for the given adsorption configuration is subsequently defined as,
$\Delta\epsilon_{\mathrm{ad}}=\epsilon_{\mathrm{tot}}-(\epsilon_{X}+\epsilon_{\mathrm{sub}})\ .$ (1)
In cases of adsorption, such convention yields $\Delta\epsilon_{\mathrm{ad}}<0$. It should be noted that this adsorption energy is derived from the $0~{\rm K}$ potential energy surface (PES) without thermodynamic corrections; thus, the results must be adjusted for thermodynamic effects and zero-point energy before application in finite-temperature physical scenarios (Appendix A).
II.1 Adsorption onto Carbonaceous Surfaces
Carbonaceous dust grains are considered to exist in graphitic-like states, including graphites, oxidized graphites, graphene, and even polycyclic aromatic hydrocarbons (PAHs) at the smallest sizes $\lesssim 0.01~{\rm\mu m}$ (see also Draine 2011), or as amorphous carbon. In terms of their astrophysical properties, these materials are observed to behave similarly, with the surface layer of carbon atoms dominating due to the weak vdW coupling between graphitic layers.
II.1.1 Single-molecule Adsorption onto Graphene
The adsorption of single molecules onto a single layer of graphene and amorphous carbon is first investigated as fiducial cases in this study. Additional calculations are conducted to verify that various situations, such as multiple adsorbates, yield quantitatively consistent results regarding the properties of interest. The construction of the graphene substrate is described in Appendix B.1.
Beginning with water molecules, following Brandenburg et al. (2019), the molecules are positioned above the center of a carbon hexagon, with orientations classified as “0-leg”, “1-leg”, and “2-leg” (illustrated in Figure 2). For each orientation, a series of $\Delta E_{\rm abs}$ values is obtained at different displacements $d$ from the substrate surface, measured from the plane of carbon nuclei to the center-of-mass of the adsorbate molecule. It should be noted that during this step, the ionic structure of the system is fixed to manually iterate through configuration space. This setup ensured that the adsorbate configuration—particularly its orientation and distance to the substrates—remained unchanged, allowing correct energy curves to be obtained. The resulting energy curves for these configurations are presented in Figure 2. The difference between our approach (the r2SCAN + rVV10 method) and the QMC method for the 1-leg case (Ma et al., 2011) is $\lesssim 0.05~{\rm eV}$, well within the range of variations among different XC methods (Hamada, 2012). Similar calculations for $\mathrm{H_{2}}$ and $\mathrm{CO}$ are performed following the same paradigm, with molecular orientation notations illustrated in Figure 2. For completeness and verification purposes, the adsorption of helium is also examined, yielding an adsorption energy of $|\Delta\epsilon_{\mathrm{ad}}|\lesssim 0.07~{\rm eV}$.
Figure 3: Example of $\mathrm{H_{2}O}$ single-molecule adsorption onto carbonaceous surface under various situations, placed above: (a) “doped” oxygen atom, (b) doped nitrogen atom, (c) hydrogen confined in the bulk substrate, and (d) surface not covered by hydrogen.
Table 1: Physical parameters of adsorbates on amorphous carbon grain surfaces.
Adsorbates Neighbors $\nu/{\rm THz}^{\dagger}$ $\xi/{\rm THz}^{\dagger}$ $|\Delta\epsilon_{\mathrm{ad}}|/{\rm eV}$ with neighbors
0 11 $\mathrm{max}^{\ddagger}$
$\mathrm{H_{2}O}$ $\mathrm{H_{2}O}$ $\sim 3.5$ $\sim 2.0$ 0.16 0.36 0.87
$\mathrm{CO}$ 0.26 0.35
$\mathrm{CO}$ $\mathrm{CO}$ $\sim 3.6$ $\sim 3.0$ 0.10 0.13 0.14
$\mathrm{H_{2}O}$ 0.24 0.27
$\mathrm{H_{2}}$ $\mathrm{H_{2}}$ $\sim 5.0$ $\sim 3.5$ 0.028 0.029 0.029
Note. — $\dagger$: The eigen frequencies of vertical ($\nu$) and horizontal ($\xi$) vibrations of adsorbates on the adsorption sites, estimated by finite difference derivatives.
$\ddagger$: The maximum number of neighbors on the surface is generally 4 nearest and 4 second-nearest.
*: Interactions between $\mathrm{H_{2}}$ and other types of adsorbates are similarly weak and not presented.
II.1.2 Adsorption onto amorphous carbon surfaces
The realistic composition of interstellar and protostellar carbonaceous grains is considered to differ considerably from graphites or graphene. According to experimental studies Jager et al. (1998), these grains are more likely to be amorphous carbon, which in various cases may be doped with oxygen and hydrogen at approximately $\sim 1~\%$ of atomic concentration. Hydrogen atoms are expected to be located predominantly at grain surfaces due to their inability to achieve coordination numbers greater than unity. Oxygen atoms may exist in both the bulk substrate and near the surface; however, numerical experiments indicate that surface oxygen atoms are always saturated by capturing hydrogen atoms because of intense dipole moments. In practice, while the cases with doped atoms are also calculated, only hydrogen and carbon atoms are elaborated and explored in details. The construction of amorphous carbon substrates is explained in Appendix B.2.
Numerical experiments for single-molecule adsorption energy are conducted above the amorphous carbon substrate. The adsorbate is positioned at five different horizontal locations above the surface, with offsets from the horizontal center of the substrate of transverse size $(9.09~\text{\AA })^{2}$, as defined by:
$\begin{split}(\Delta x,\ \Delta y)&\in\{(-1,-1),(-1,1),(1,-1),(1,1),(0,0)\}\\ &\times\left(\dfrac{9.09~\text{\AA }}{4},\dfrac{9.09~\text{\AA }}{4}\right)\ .\end{split}$ (2)
At each site, ionic structures are relaxed to locate the energy minimum before the adsorbate is moved sufficiently far from the substrate to evaluate $\Delta\epsilon_{\mathrm{ad}}$ using equation (1). Single-molecule adsorption energies in Table 1 are obtained by averaging $\Delta\epsilon_{\mathrm{ad}}$ over the five representative sampling sites. It is acknowledged that the inherent stochasticity and structural heterogeneity of amorphous carbon surfaces could be better represented by a larger ensemble of sampling points characterizing such disordered media. Nonetheless, an exhaustive sampling of the near-infinite configuration space of local adsorption sites is computationally prohibitive. The amorphous carbon substrate model generated for this study (Appendix B.2) also exhibits some degree of structural homogeneity within the supercell that facilitates representative sampling of the dominant binding environments. In fact, supplementary DFT calculations have also been performed on a wider set of prospective sites and alternative amorphous carbon realizations (not shown in this paper), whose consequent adsorption energies consistently fall within the range established by the five primary sampling points (except a few extreme cases, where the adsorbates eventually reside in the sites whose $|\Delta\epsilon_{\rm ad}|$ still stay in the consistent range after structure relaxation).
In It is noted that replacing approximately $\sim 2~\%$ of carbon atoms with oxygen yielded nearly identical results, as oxygen atoms are rarely exposed directly (due to hydrogen surface coverage) and thus contributed minimally to interactions with adsorbates. Through deliberate exposure of oxygen and nitrogen atoms at the surface (e.g., Figure 3), the adsorption energy of $\mathrm{H_{2}O}$ via hydrogen bonds is measured to be $0.35~{\rm eV}$ (above oxygen) and $0.29~{\rm eV}$ (above nitrogen), respectively. The adsorption energy of $\mathrm{CO}$ near doping atoms exhibited no significant variations.
When substrate surface coverage fractions become non-negligible, interactions between adsorbates may play a crucial role in modifying their kinetic behaviors. The vdW interactions and hydrogen bonds between adjacent molecules increase the barriers for desorption and hopping, with intensity dependent on both the numbers and types of neighbors. These effects are quantified through numerical experiments based on amorphous carbon substrates. The interaction energy is measured through the following procedures:
  1. 1.
    Initial site pre-optimization: The “primary” adsorbate of interest is first placed and relaxed at a previously descreibed single-molecule adsorption site via full geometry optimization.
  2. 2.
    Initial neighbor positioning: Neighboring molecules are introduced at radial distances approximating the characteristic intermolecular separations found in their respective bulk solid phases (e.g., the hydrogen-bonding network of amorphous solid water when the primary molecule and its neighbors are all $\mathrm{H_{2}O}$). The azimuthal angles of these neighbors are selected so that the neighbors are close to their likely single-molecule adsorption site if possible (e.g., the nearest available hexagonal centers or “hollow” sites over the amorphous surface), or in random otherwise (while the separations between neighbors are still guaranteed).
  3. 3.
    Constrained adsorbate relaxation: DFT structural optimization is conducted where the degrees of freedom of all adsorbed molecules are relaxed to a local energy minimum. During this step, the coordinates of the substrate atoms are held fixed to maintain the pre-equilibrated surface morphology.
  4. 4.
    Adsorption energy measurement: Adsorbate of interest is moved sufficiently far from the substrate ($\Delta z\gtrsim 10$$\text{\AA }$), and configurations of all remaining atoms are relaxed to re-evaluate the total energy.
Energy differences measured through these steps yielded the “neighbor-enhanced” total binding energy. For tests involving adsorbates with one neighbor, adsorbate pairs are positioned at five different sites (with centers of mass offsets following equation 2) to obtain average enhanced binding energies. The overall binding energy (including adsorption) is observed to increase sub-linearly, as vdW forces exhibit saturation behavior and hydrogen bonds require specific molecular orientations. Tests are also conducted to determine the numbers of different neighbors that maximize absolute binding energy values, beyond which binding energy decreases due to repulsive forces. Similar to the single-molecule cases, depleting the possibilities over the heterogeneous amorphous carbon surfaces is virtually not practical. Nonetheless, because of the qualitative homogeneity of the model amorphous surface, it is expected that such procedures could still reasonably sample the typical situations that are relevant in the following calculations.
II.2 Adsorption onto Silicates
Figure 4: Configurations of adsorbed $\mathrm{CO}$ molecules (upper row) and $\mathrm{H_{2}O}$ molecules (lower row) over different crystal surfaces (indicated above each column) of the model silicate $\mathrm{MgSiO_{3}}$. It should be noted that adsorption of $\mathrm{H_{2}O}$ over the $(100)$ surface results in molecular dissociation.
In comparison to carbonaceous grains, silicate compositions are recognized to be considerably more complex. To maintain model tractability while preserving physical relevance, iron-poor enstatite $\mathrm{MgSiO_{3}}$ (Pnma space group using the Materials Project data, see also Jain et al. 2013) is selected as a proxy for magnesium silicate substrates. This choice is motivated by the abundance of magnesium, which among alkali and alkaline earth metals in the ISM and PPDs, typically matches silicon abundance levels.The details of silicate substrate model construction are explained in Appendix B.3.
Admittedly, the model crystalline Pnma enstatite is unable to fully represent the amorphous enstatite, a commonly encountered phase of astrosilicates also in PPDs (see Henning, 2010, and references therein). Practical representation of disordered systems in the ab-initio DFT frameworks, nevertheless, typically involves significant computational trade-offs. While crystalline substrates can be accurately modeled with relatively compact supercells ($\sim 10^{2}$ atoms), a physically realistic amorphous structure requires substantially larger, disordered models (typically $\sim 10^{3}$ atoms) to minimize periodic artifacts, making systematic potential energy surface mapping with advanced metaGGA and van der Waals functionals (e.g., r2SCAN+rVV10 adopted in this work) prohibitively expensive.
In this study, crystalline enstatite is adopted as a generic proxy for its amorphous counterpart, becaus (1) both phases share very similar stoichiometry, providing a chemically representative building block for magnesium-rich silicates; and (2) surface reactivity is primarily governed by local coordination environments rather than long-range order, with adsorption dominated by undercoordinated Mg and Si sites whose chemical interactions remain consistent across phases. To better approximate the morphological and chemical heterogeneity of a disordered surface within these computational constraints, calculations were performed across multiple distinct facets including (100), (010), and $(001)^{\pm}$, which expose varying ratios of Mg and Si atoms in diverse coordination states, effectively bracketing the range of local environments encountered on a realistic silicate surface.
The specific atomic composition of silicate substrates beneath adsorbate molecules is found to determine adsorption properties. Initial investigations positioned molecules with varying orientations at different heights above different atomic sites. After identifying configurations with approximately minimal energy, molecules are placed accordingly on the substrate, and ionic structures are evolved to achieve relaxation, eliminating total stress to the desired accuracy ($\lesssim 10^{-2}~{\rm eV}~\text{\AA }^{-1}$) and thereby locating energy extrema. In astrophysical conditions relevant to PPDs, it is possible that the silicate surfaces could be covered by chemisorbed hydrogen. Additional tests were performed on surfaces partially covered with these hydrogen (not shown in this paper), which preferentially saturates exposed under-coordinated oxygen atoms while leaving some Mg and Si coordination centers accessible. These tests indicate that hydrogen coverage does not significantly alter the adsorption energies of $\mathrm{H_{2}O}$ and CO on enstatite surfaces (with changes limited to $\lesssim 0.2~{\rm eV}$), which is of secondary importance compared to the total chemisorption energy of $|\Delta\epsilon_{\rm ad}|\gtrsim 1~{\rm eV}$. This is consistent with expectation, as non-closest-neighbor conditions do not strongly affect covalent coordinate bonds.
Configurations converged at energy extrema are illustrated in Figure 4, with corresponding adsorption energies presented in Table 2. The adsorption energy values for $\mathrm{H_{2}O}$ and $\mathrm{CO}$ are generally observed to be one order of magnitude greater than those on carbonaceous surfaces, clearly indicating chemisorption processes. This enhancement is attributed to the coordination numbers of metal centers (Si atoms on $(001)^{-}$ and Mg atoms elsewhere) being significantly less than the saturation value of 6 in bulk $\mathrm{MgSiO_{3}}$, thereby permitting coordination bond formation and elevating adsorption energies to approximately $\sim 1~{\rm eV}$. A notable exception is identified for $\mathrm{H_{2}O}$ adsorption on the (100) surface where the molecule adsorption eventuall leads to dissociation. On this facet, exposed magnesium atoms exhibit a preferred directionality for coordination bond formation that lies nearly tangential to the surface plane, positioning the chemisorbed water molecule close to neighboring under-coordinated silicon atoms (with coordination numbers of $\sim 4$, versus 6 in the bulk). These Si sites could accept and stabilize the transferred proton, collectively lowering the activation barrier for O–H bond cleavage. In comparison, the (010), $(001)^{+}$, and $(001)^{-}$ surfaces favor coordination geometries that extend more perpendicularly from the surface, keeping sufficient distances of the water molecule from under-coordinated atoms, while exposing more fully coordinated Si–O–Si bridging configurations that are less susceptible to protonation.
Chemical bonding is observed to constrain molecular orientations following convergence. In $\mathrm{H_{2}O}$, the oxygen atom functions as the electron donor, while in $\mathrm{CO}$, the carbon atom serves this role. Given the sufficiently high band intensities and their strong orientation dependence, numerical experiments confirmed that vdW or hydrogen bond interactions between adjacent adsorbates could be neglected on silicate substrates. The increased elasticity resulting from stronger bonds also produced higher eigen frequencies of molecular oscillation ($\nu$ and $\xi$). In contrast, the inability of $\mathrm{H_{2}}$ to form coordination bonds resulted in negligible adsorption energy.
Table 2: Physical parameters of adsorbates on different surfaces of model silicate $\mathrm{MgSiO_{3}}$.
Adsorbates $\nu/{\rm THz}$ $\xi/{\rm THz}$ $|\Delta\epsilon_{\mathrm{ad}}|/{\rm eV}$ on surfaces
$(100)$ $(010)$ $^{-} (001)+(001)^{+}$
$\mathrm{H_{2}O}$ $\sim 13$ $\sim 6.0$ -$^{**}$ $1.50$ $1.32$ $1.17$
$\mathrm{CO}$ $\sim 6.7$ $\sim 5.5$ $1.23$ $0.95$ $0.74$ $0.45$
$\mathrm{H_{2}}$ $\sim 6.3$ $\sim 3.0$ $0.015$ $0.008$ $0.016$ $0.025$
Note. — Similar to Table 1, but the neighbor conditions are not included as the molecules are chemisorbed onto silicate surfaces (see also §II.2).
**: Molecule dissociated after chemisorption.
III Kinetic Monte Carlo Simulations of Grain Surfaces
Surface coverage conditions of dust grains play a fundamental role in regulating coagulation processes through their influence on sticking behavior and surface energy dynamics. To systematically investigate these coverage conditions, we developed a specialized kinetic Monte Carlo (KMC) simulation code, optimized for GPU acceleration to handle the computational demands of these complex simulations [^1] https://github.com/wll745881210/kmc_ads. whose general procedures are explained in Appendix C. The zero-temperature results obtained in §II need thermodynamic corrections before being applied in the KMC simulations (see Appendix A). It is also noticed that the KMC approach itself inherently sacrifices the accuracy of ab-initio calculations by discretizing events and transition rates on a square lattice, rather than employing rigorous off-lattice simulations that directly utilize consistent DFT calculations. In this work, the transition rates are also approximate values derived from limited sampling of substrate conditions, based on adsorption and binding energies obtained from a limited-size amorphous carbon model or from crystalline enstatite as a proxy for amorphous silicate. Nevertheless, the purpose of conducting KMC simulations is to balance this trade-off, to obtain improved statistical sampling over larger systems at the cost of single-process accuracy. By using test results based on a model amorphous carbon surface with reasonable homogeneity, and by bracketing amorphous silicate surface environments with multiple crystalline facets, the KMC results are expected to retain reasonable semi-quantitative reliability.
III.1 Setups of KMC simulations
The KMC simulations are implemented on two-dimensional periodic square lattices designed to emulate grain surfaces, where each lattice point corresponds to a discrete adsorption site. This discretization approach allows for efficient modeling of molecular adsorption, desorption, and migration processes across the surface landscape. The KMC framework incorporated both square and hexagonal lattice geometries to assess potential structural dependencies in surface coverage simulations. Comparative analyses revealed that hexagonal lattices produced nearly identical results to square lattices in terms of surface coverage conditions, indicating that the fundamental physical processes governing molecular adsorption exhibit minimal sensitivity to lattice symmetry. Consequently, for enhanced computational efficiency and analytical clarity, only square lattice implementations are retained for detailed presentation and comparative analyses throughout this investigation.
Specific initial and terminal conditions are established to ensure physical relevance and numerical stability. Each simulation commenced with a pre-defined “condensation nucleus” consisting of a $4\times 4$ patch of water molecules positioned on the surface. This initialization strategy prevents numerical artifacts associated with supercooling or superheating conditions that could yield unreliable results. Simulation convergence is defined by either the complete disappearance of the condensation nucleus or the achievement of surface coverage exceeding $99\%$. In cases of extensive coverage, where second and higher molecular layers begin to form, analyses are deliberately restricted to the first monolayer to maintain interpretational clarity.
III.2 Surface coverage conditions simulated by KMC
The divergent surface coating behaviors observed in protoplanetary disks (PPDs) directly originate from the fundamentally different adsorption physics governing carbonaceous and silicate grains. Within a representative PPD model approximating early Solar System conditions, we conduct kinetic Monte Carlo simulations based on vdW-DFT computational results. These KMC implementations comprehensively account for factors influencing desorption rates, including finite-temperature thermodynamic corrections and inter-adsorbate interactions parameterized through the eigen frequencies of adsorbate vibrations (Tables 1 and 2).
Simulations are conducted for dust grains situated in the mid-plane of a model PPD, employing density and temperature profiles consistent with a young stellar object of bolometric luminosity $L\sim 4L_{\odot}$. The temperature profile follows $T\propto R^{-0.4}$ due to surface flaring effects Chiang and Goldreich (1997), with specific forms given by:
$\begin{split}&T\simeq 280~{\rm K}\left(\dfrac{R}{\mathrm{AU}}\right)^{-0.4},\\ &\rho_{0}\simeq 3.8\times 10^{-10}~{\rm g}~{\rm cm}^{-3}\left(\dfrac{R}{\mathrm{AU}}\right)^{-2.8}.\end{split}$ (3)
This mid-plane density corresponds to a surface density profile of $\Sigma\simeq 600~{\rm g}~{\rm cm}^{-1.5}(R/\mathrm{AU})^{-1.5}$. Molecular number densities are derived assuming standard abundance ratios: $X(\mathrm{H_{2}})\simeq 1$, $X(\mathrm{CO})\simeq 1.8\times 10^{-4}$, and $X(\mathrm{H_{2}O})\simeq 1.4\times 10^{-4}$. Helium atoms are excluded from consideration due to their negligible adsorption efficiency on grain surfaces.
[…]
Across the radial domain of $0.1<(R/\mathrm{AU})<200$, KMC simulations revealed distinct surface coverage patterns for carbonaceous surfaces and the $(001)^{-}$ silicate surface, as illustrated in Figures 5. Silicate surfaces undergo a transition from bare to coated states at $R\lesssim 0.2~\mathrm{AU}$, a consequence of the substantial chemisorption energies characteristic of $\mathrm{CO}$ and $\mathrm{H_{2}O}$ on these substrates. In contrast, KMC results clearly identify a snowline on grain surfaces for water on carbonaceous grains at $R\simeq 8~\mathrm{AU}$ ($T\simeq 120~{\rm K}$), interior to which surfaces remain uncoated due to elevated thermal energies (Figure 5). This snowline represents a distinct concept from the conventional snowline marking $\mathrm{H_{2}O}$ ice condensation at $T\sim 170~{\rm K}$ (Sasselov and Lecar, 2000), with the former being more directly relevant to substrate material coagulation processes. Employing the previously assumed adsorption energy of $\simeq 0.45~{\rm eV}$ would reposition the $\mathrm{H_{2}O}$ snowline at $R\simeq 2~\mathrm{AU}$.
The fundamental differences in adsorption behavior between carbonaceous and silicate substrates have likely remained obscured in previous research due to systematic misinterpretations of temperature-programmed desorption (TPD) experimental data. This misinterpretation stems from the complex nature of water cluster formation on substrate surfaces. Within water molecule clusters, each molecule typically participates in two hydrogen bonds on average, mirroring the structural configuration of hexagonal ice crystals (Ih). This hydrogen bonding network contributes an internal binding energy of approximately $\epsilon_{\mathrm{hb}}\sim 0.42~\mathrm{eV}$ per molecule (where the subscript “hb” denotes hydrogen bonds; see also Silverstein et al. (2000)).
Conventional TPD experiments are conducted under conditions substantially different from astrophysical environments: temperatures well below the triple point of water and gas-phase molecule number densities orders of magnitude higher than those found in protoplanetary disks or the interstellar medium. These experimental conditions promote the formation of extensive water clusters on substrate surfaces Leenaerts et al. (2009); Ma et al. (2011); Björneholm et al. (2016). Consequently, molecular addition or removal processes during TPD measurements primarily occur at the interfaces of these water clusters rather than at the actual substrate-adsorbate boundary. This clustering phenomenon could lead to a possible misinterpretation of TPD data: the slopes observed in Arrhenius plots actually reflect $\epsilon_{\mathrm{hb}}/k_{\mathrm{B}}$, the energy scale associated with hydrogen bond breaking, rather than the true substrate adsorption energy $\Delta\epsilon_{\mathrm{ad}}/k_{\mathrm{B}}$. For silicate substrates, the situation is further complicated by the exceptional strength of chemisorption bonds, which remain intact at typical protoplanetary disk temperatures ($T\lesssim 10^{3}~{\rm K}$). Under these conditions, any molecular desorption processes are governed by hydrogen bond rupture within water clusters, again causing TPD measurements to yield $\epsilon_{\mathrm{hb}}$ rather than the fundamental substrate adsorption energy. Accurate interpretation of adsorption mechanisms and binding energies is therefore essential for reliable modeling of astrophysical processes involving dust grains.
IV Discussions
We have combined vdW-corrected ab-initio DFT calculations and kinetic Monte Carlo simulations to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces, and how these material-dependent adsorption energies, combined with KMC simulations, translate into different surface coverage patterns. The results reveal a clear dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicates, with important consequences for grain growth and volatile partitioning. We notice that, although crystalline enstatite is employed as a proxy for more complex amorphous silicates, the fact that chemisorption energies on silicate surfaces are an order of magnitude stronger than on carbonaceous surfaces remains relatively robust: the adsorption energy magnitudes ($|\Delta\epsilon_{\rm ad}|\gtrsim 1~{\rm eV}$ for $\mathrm{H_{2}O}$ and $\sim 0.5-1~{\rm eV}$ for CO) are not significantly influenced by the crystalline nature of the substrate (Tables 1 and 2).
[…]
IV.1 Dependence of snowline locations on
evolutionary history
The initial surface conditions of dust grains could be crucial in determining the locations and characteristics of molecular snowlines in protoplanetary disks. When grain surfaces begin with complete $\mathrm{H_{2}O}$ molecular coatings rather than initially bare configurations, the adsorbed water molecules demonstrate significantly enhanced thermal stability due to pre-existing intermolecular binding networks. This phenomenon creates metastable surface conditions that share fundamental physical characteristics with superheating effects observed in other first-order phase transitions.
The enhanced thermal stability arises from the collective binding energy provided by hydrogen bonding networks between adjacent water molecules. These networks create energy barriers that substantially exceed the single-molecule adsorption energy, allowing molecular layers to persist at temperatures far above those predicted by simple adsorption models. Consequently, the ultimate surface coverage conditions and associated gas-phase condensation processes (including potential molecular depletion) become strongly dependent on the specific thermal and dynamical evolution history experienced by dust grains.
The divergence between different evolutionary pathways can be dramatic. When pre-hydrated grains undergo gradual heating from cold initial conditions, they may maintain surface coatings up to temperatures approximately a few tens of Kelvin (typically $\sim 20-30~{\rm K}$) higher than the equivalent snowline temperature for condensation onto initially bare surfaces (not shown in this paper). This effect arises because the formation of new ice layers requires nucleation events that possess significant activation energies, while the removal of existing layers requires overcoming the collective binding energy of the entire molecular network. Therefore, accurate and consistent determination of snowline locations and condensation conditions necessitates comprehensive modeling that incorporates the full thermodynamic history of disk material, including radial drift trajectories and temporal temperature variations. The assumption of instantaneous equilibrium between adsorption and desorption processes proves insufficient for capturing these history-dependent effects, requiring instead the implementation of kinetic models that track the evolutionary pathways of individual grain populations throughout their disk residence.
The final surface coverage of CO on dust grains is not a static property but depends on the specific thermal and dynamical evolution path of the dust population. The initial conditions of grains, such as starting as “dry grains” that are too cold to have undergone significant thermal processing, can preset the availability of binding sites and the efficiency of subsequent ice mantle formation. Furthermore, the radial transport of grains plays a crucial role. Icy grains drifting inward from the outer disk sublimate upon crossing their respective snowlines, temporarily enhancing the local gas-phase abundance before potentially re-condensing onto other grain surfaces under the new thermal conditions.
Consequently, history-dependent snowline locations could vary in different disks, occurring at radii corresponding to different temperatures. Simply inferring their locations based on a fixed temperature (e.g., $20~{\rm K}$ for CO) may therefore lead to mismatches. Dust rings and gaps have been found to be nearly ubiquitous in disks. The prevalent hypothesis for their origin is that they are caused by planet–disk interactions (Dong et al., 2015, 2017; Dong and Fung, 2017), while the snowline hypothesis (Zhang et al., 2015) has been questioned due to mismatches between the observed ring/gap locations and the expected snowline locations (Long et al., 2018; Bae et al., 2018; van der Marel et al., 2019). Re-examining the connection between the two, while accounting for each disk’s thermal and dynamical history, is therefore needed.
IV.2 Impact of multi-species adsorption on CO surface snowlines and gas depletion
The analysis of surface coverage conditions reveals that, based on the disk temperature and hydrodynamic profiles, $R\sim 20~\mathrm{AU}$ approximately divides the grain surface snowline for $\mathrm{CO}$ on amorphous carbon grains, corresponding to a local temperature of $T\simeq 80~{\rm K}$ according to the disk temperature profile defined in equation (3). This temperature appears anomalously high when compared to existing experimental and computational studies, which typically report CO condensation temperatures around $\sim 20~{\rm K}$ (Cazaux et al., 2017). However, detailed examination of Figure 5 indicates that CO condensation initiation is not an isolated process but rather occurs through coordinated condensation with $\mathrm{H_{2}O}$ molecules.
The underlying mechanism for this elevated condensation temperature becomes evident upon inspection of Table 1, which demonstrates that the interaction energy between a $\mathrm{CO}$ molecule and an adjacent $\mathrm{H_{2}O}$ molecule significantly exceeds that between two $\mathrm{CO}$ molecules. This enhanced intermolecular binding substantially increases the effective adsorption energy of CO molecules when they are incorporated within a water ice matrix. The resulting configuration resembles “cocrystal” structures, where CO molecules become trapped within the lattice of water ice (distinct from traditional clathrate hydrates, particularly in their interaction characteristics with $\mathrm{H_{2}}$). This cocrystal formation enables the stabilization of surface-phase CO molecules at temperatures far exceeding their typical triple point, highlighting the critical importance of considering multi-species interactions when modeling gas-phase depletion processes in protoplanetary disks.
It should be further noted that while the primary investigations in this work focus on single-layer adsorption phenomena, additional studies of second-layer $\mathrm{H_{2}O}$ adsorption on silicate-supported water monolayers revealed a single-molecule adsorption energy of $0.28~{\rm eV}$, resulting from the combined effects of hydrogen bonding and van der Waals interactions. When neighbor interactions are incorporated, this corresponds to a “multiple-layer snowline” at $\sim 1~\mathrm{AU}$ as indicated by supplementary KMC simulations (not directly illustrated in this work), a finding consistent with conventional snowline scenarios.
Nevertheless, the single-layer adsorption conditions remain particularly relevant for several reasons. First, grain coagulation and subsequent planet formation processes are predominantly governed by single-layer surface conditions, as the sticking versus bouncing outcomes during collisions are primarily determined by center-of-mass collision energies that are most sensitive to the last adsorbed molecular layer. Second, the fundamental physical mechanism enabling CO condensation at elevated temperatures through enhanced binding with $\mathrm{H_{2}O}$ neighbors remains equally valid in multiple-layer scenarios, as the crucial factor is the strengthened binding energy and consequent confinement of $\mathrm{CO}$ molecules within the water ice matrix, regardless of the total number of adsorbed layers.
IV.3 Molecule Condensation and Gas-Phase Depletion
Our calculations indicate that the CO midplane snowline lies significantly closer to the central star than conventional estimates suggest. In a representative PPD around a solar-type star with $T(1~\mathrm{AU})\simeq 150~{\rm K}$, the CO snowline is expected at around $30~\mathrm{AU}$ (e.g., Öberg et al., 2023). When adopting the higher midplane temperature used in our disk model ($280~{\rm K}$ instead of $150~{\rm K}$ at $1~\mathrm{AU}$), this discrepancy becomes even more pronounced. This inward shift of CO snowline in our model mainly arises from enhanced $\mathrm{CO}$$\mathrm{H_{2}O}$ binding, which promotes earlier condensation of CO onto water-rich grain surfaces.
Observationally, molecular snowline locations can be inferred from chemical tracers of gas-phase abundances, such as $\mathrm{N_{2}H^{+}}$ for CO, which usually place the CO snowline beyond 40 AU in disks around solar-type stars (Qi et al., 2013, 2019). However, these diagnostics are not always uniquely linked to volatile depletion at the corresponding snowline and can be affected by optical depth effects, non-thermal desorption, or chemical processing (van ’t Hoff et al., 2017). These measurements typically trace the traditional $20~{\rm K}$ condensation front, where the CO gas abundance drops dramatically. In contrast, our models predict the onset of CO condensation significantly closer to the star, at a region not directly captured by such observations. This suggests that a substantial disk region may contain partially CO-ice-coated grains, effectively extending the CO-ice zone inward compared to classical estimates.
[…]
Our models also predict that $\mathrm{H_{2}O}$ ice can sublimate at lower temperatures and thus at larger disk radii, when it is coated on carbonaceous grains. These shifts of snowline locations have important implications for interpreting the inner disk chemistry revealed by JWST MIR spectra. The initial gas-phase C/O ratio in disks is largely determined by the relative positions of major volatile snowlines. For example, it is expected to be lower inside the $\mathrm{H_{2}O}$ snowline than in the region between $\mathrm{H_{2}O}$ and $\mathrm{CO}$ snowlines (Öberg et al., 2011). Current interpretations of JWST spectra often invoke the sublimation of icy grains as they drift inward and cross various snowlines (e.g., Mah et al., 2023). Recent models that include CO entrapment in water ice suggest elevated C/O and C/H ratio inside the water snowline, by up to a factor of 10 over $1~{\rm Myr}$ (Williams et al., 2025). This process is analogous to the closer-in onset of CO condensation predicted in our calculations. Precise knowledge of snowline locations and their temperature dependencies is thus essential for modeling disk chemistry. The outward shift of the water snowline predicted here adds an additional layer of complexity that future chemo-dynamical models should consider. Moreover, the lower sublimation temperature of water ice implies that $\mathrm{H_{2}O}$ vapor may persist at colder temperatures, motivating future far-infrared observations with facilities such as PRIMA to probe water beyond the canonical snowline.
V Summary
We have combined vdW-corrected ab-initio DFT calculations (§II) and kinetic Monte Carlo simulations (§III) to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces. Although complete representations of amorphous carbon and amorphous silicates are computationally prohibitive, the surfaces of model amorphous carbon and crystalline enstatite have been sampled as proxies. The strong chemisorption energy ($\gtrsim 1~{\rm eV}$) of $\mathrm{H_{2}O}$ and CO molecules on silicate surfaces renders the prospective error from using such proxies relatively less important. By sampling various facets of crystalline enstatite, the bracketing approach yields results for silicate surfaces that are reasonable, semi-quantitative approximations to realistic amorphous silicates.
The material-dependent adsorption energies, when incorporated into the KMC framework, lead to distinct surface coverage patterns and well-defined snowlines on grain surfaces in a representative protoplanetary disk model (Figure 5). Our calculations have also demonstrated a dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicate grains (Tables 1 and 2), which results in carbonaceous grains losing their volatile coatings in the inner disk while silicate grains remain molecularly coated over a much wider radial range.
Building on this physical picture, we show that history-dependent initial surface conditions and evolutionary paths can shift snowline locations (§IV.1), that multi-species adsorption and enhanced $\mathrm{CO}$ binding in $\mathrm{H_{2}O}$-rich ices can move the effective CO surface snowline inward and strengthen gas-phase CO depletion (§IV.2), and that these effects together impact disk gas-mass estimates, the interpretation of snowline-sensitive tracers, and JWST spectra in planet-forming disks (§IV.3).
From an observational perspective, the carbonaceous snowline on carbonaceous grain surfaces developed in this work should influence carbon abundance patterns in extrasolar planetary systems, as the lack of “glue molecules” on carbonaceous grains at warm temperatures can lead to the generic scarcity of carbon in regions closer to the host star. Future observational campaigns, particularly those utilizing improved characterization of elemental abundances in exoplanetary atmospheres, will provide crucial tests of these predictions across a broader range of planetary system architectures. Such data may additionally offer constraints on potential planetary migration histories. This work establishes a foundation for investigating key astrophysical processes through a modernized approach grounded in the fundamental physics of molecules, atoms, and electrons, bridging microscopic interactions with macroscopic astrophysical phenomena.
This work is supported by the NSFC General Project 12573067. This research was enabled partially by the High-performance Computing Platform of Peking University. We thank our collegues: Di Li, Luis C. Ho, Jeremy Goodman, Ahmad Nemer, Satoshi Okuzumi, Kengo Tomida, Kazuyuki Omukai, Mordecai Mac-Low, Xiao Hu, Siyi Feng, for helpful discussions and comments.
Appendix A Thermodynamic Corrections for Adsorption Energies
The thermodynamic treatment of adsorbate-substrate systems requires careful consideration of their inherent complexity compared to isolated molecular systems. Dust grain substrates possess substantially larger heat capacities than individual adsorbate molecules and maintain stiff thermal equilibrium with the diffuse radiation fields characteristic of interstellar media and protoplanetary disks Chiang and Goldreich (1997). Consequently, these substrates effectively function as isothermal heat reservoirs during adsorption-desorption processes. When molecular desorption occurs, the thermal relaxation timescale in the gas phase through collisions with other molecules can be estimated following Draine (2011):
$\begin{split}\tau_{\mathrm{coll,gas}}&\equiv[n_{\mathrm{H_{2}}}\langle\sigma v_{\mathrm{th}}\rangle]^{-1}\\ &\sim 1~\mathrm{s}\times\left(\dfrac{T}{10^{2}~{\rm K}}\right)^{-1/2}\left(\dfrac{\rho_{\mathrm{gas}}}{10^{-14}~{\rm g}~{\rm cm}^{-3}}\right)^{-1/2}.\end{split}$ (A1)
Under typical protoplanetary disk mid-plane conditions, $\tau_{\mathrm{coll,gas}}$ exceeds the desorption timescale by more than ten orders of magnitude, effectively halting thermal energy exchange between desorbed molecules and the surrounding gas phase. This dramatic timescale separation necessitates distinct thermodynamic treatments for different molecular degrees of freedom: molecular motion perpendicular to the substrate surface (the desorption direction) should be treated as adiabatic, with the internal energy change $\Delta\epsilon_{\mathrm{ad}}$ serving as the relevant thermodynamic potential. Conversely, motion parallel to the surface maintains continuous thermal contact with the substrate heat reservoir, requiring that interactions with neighboring adsorbates during hopping and desorption processes be treated isothermally.
The vibrational characteristics of adsorbed molecules were determined through analysis of their quasi-harmonic oscillations near energy minima. Following convergence to minimal energy configurations with fixed substrate ionic structures, the VASP software package was employed to evaluate second derivatives of the potential energy surface via finite differences. Diagonalization of the resulting Hessian matrices yielded eigenfrequencies for adsorbate oscillation modes, enabling precise thermodynamic corrections.
For the adiabatic dimension (perpendicular motion), thermodynamic corrections primarily account for the transition from free translational motion to confined quasi-harmonic oscillations. The internal energy correction incorporating zero-point energy (ZPE) contributions is given by:
$\delta\epsilon_{\perp}=h\nu\left(\dfrac{1}{2}+\dfrac{1}{e^{\beta h\nu}-1}\right)-\dfrac{1}{2}k_{\mathrm{B}}T\ ,$ (A2)
where $\nu$ represents the vertical eigenfrequencies tabulated in Tables 1 and 2, and $\beta\equiv(k_{\mathrm{B}}T)^{-1}$.
For the isothermal dimensions (parallel motion), the Helmholtz free energy serves as the appropriate thermodynamic potential. The binding energy correction is derived from the free energy difference between the bounded oscillatory state and the unbounded translational state:
$\begin{split}&f_{\mathrm{osci}}=\dfrac{h\xi}{2}+\dfrac{1}{\beta}\ln\left[1-e^{-\beta h\xi}\right]\ ,\\ &f_{\mathrm{trans}}=-\dfrac{2}{\beta}\ln\left[\delta l\left(\dfrac{h^{2}\beta}{2\pi\mu}\right)^{-1/2}\right]\ ,\end{split}$ (A3)
where $\xi$ denotes the horizontal eigenfrequencies from Tables 1 and 2, $\mu$ represents the molecular mass, and $\lambda_{\mathrm{th}}\equiv(h^{2}\beta/2\pi\mu)^{1/2}$ corresponds to the thermal de Broglie wavelength. The parameter $\delta l$ describes the average intermolecular distance on the substrate surface, which varies with both gas-phase and adsorbed-phase conditions. For computational simplicity, $\delta l\sim 10$$\mathrm{\AA }$ was adopted, approximately corresponding to the distance where interactions between adjacent adsorbates become negligible. The logarithmic dependence on $\delta l$ ensures that the results remain insensitive to its exact value, requiring only order-of-magnitude accuracy. Increasing $\delta l$ elevates the value of $\delta f_{\parallel}$, thereby generally weakening the binding between adsorbates.
Appendix B Construction of Dust Grain Substrate Models
B.1 Graphene Substrate Construction
[…]
Brillouin zone sampling was conducted using a $\Gamma$-centered $1\times 1\times 1$ k-point grid during initial DFT calculations. The adequacy of this sampling scheme was verified through comparative tests employing denser $3\times 3\times 1$ and $5\times 5\times 1$ $k$-point meshes at identified energy minima, confirming numerical reliability and convergence of the computed adsorption energies. Following initial energy curve determination, full structural relaxation of adsorbate configurations was performed to locate the true energy minima and obtain accurate minimal adsorption energies.
B.2 Amorphous Carbon Substrate Construction
A representative model for interstellar amorphous carbon grains was constructed through a multi-step computational procedure. Initial configuration generation involved placement of 75 carbon atoms on a randomly perturbed $5\times 5\times 3$ lattice within a periodic boundary condition unit cell. To emulate the hydrogenation conditions expected in interstellar environments, 30 hydrogen atoms were randomly distributed atop the carbon bulk, preferentially associating with carbon atoms exhibiting low coordination numbers.
Structural optimization was performed using DFT with full consideration of electronic interactions, allowing both ionic positions and supercell dimensions to vary during relaxation. Convergence was achieved when the total force on all atoms fell below the threshold of $10^{-2}~{\rm eV}~\text{\AA }^{-1}$. Hydrogen atoms failing to form chemical bonds with carbon during this process were subsequently removed from the system, simulating the natural saturation processes occurring in astrophysical environments.
The final optimized structure possessed a supercell with transverse dimensions of $(9.09~\text{\AA })^{2}$ and exhibited a mixed hybridization state distribution: 47% of carbon atoms in $sp^{2}$ configuration and 53% in $sp^{3}$ configuration. The resulting carbon bulk demonstrated an average thickness of $\sim 5.1$$\text{\AA }$ and a density of $\sim 2.4~{\rm g}~{\rm cm}^{-3}$, values consistent with experimentally characterized amorphous carbon materials. Surface hydrogen coverage reached $\sim 0.16~{\rm cm}^{-2}$, corresponding to approximately 1% atomic hydrogen concentration for a spherical grain of radius $0.05~{\rm\mu m}$, in excellent agreement with experimental measurements Jager et al. (1998).
B.3 Silicate Substrate Construction
Model silicate substrates were constructed using iron-poor enstatite $\mathrm{MgSiO_{3}}$ in the Pnma space group as a representative magnesium silicate system. Surface selection presented significant complexity due to the theoretically infinite variety of possible crystal facet orientations. To ensure computational tractability while maintaining physical relevance, surfaces with the lowest Miller indices, (100), (010), and (001), were selected as representative substrates. Surface terminations were specifically chosen to minimize dipole moments in the near-surface region, enhancing numerical stability and physical realism.
The (001) surface orientation exhibited particular complexity, manifesting two distinct termination types: the $(001)^{+}$ surface exposing magnesium atoms and the $(001)^{-}$ surface exposing silicon atoms. This directional asymmetry resulted from the non-centrosymmetric nature of the Pnma crystal structure. In contrast, the (100) and (010) surfaces did not exhibit such directional dependence due to inherent symmetry properties.
[…]
Appendix C General procedures of KMC simulations
KMC simulations have been performed extensively for various studies for surface physical and chemical properties (e.g. Andersen et al., 2019). Three distinct processes are modeled to occur on the simulated grain surfaces, each characterized by specific rate equations:
Adsorption. Each unoccupied surface site experiences molecular adsorption at a characteristic rate defined by $\zeta_{i}\equiv n_{i}v_{{\rm th},i}\Sigma$, representing the effective incoming flux frequency for molecular species $i$. This formulation incorporates the gas-phase number density $n_{i}$, the thermal velocity $v_{{\rm th},i}$ derived from Maxwell-Boltzmann statistics, and $\Sigma$, which denotes the effective adsorption cross-section with intrinsic sticking probability implicitly included. Through systematic analysis of solid-phase structures derived from vdW-DFT calculations, an optimal value of $\Sigma\simeq(3~\text{\AA })^{2}$ is determined, providing a consistent approximation applicable across all substrate-adsorbate combinations examined in this study.
Hopping. Molecular migration across surface sites is modeled through a hopping mechanism, where an adsorbate occupying a site with $N$ adjacent vacancies exhibits a hopping rate expressed as:
$k_{\rm hop}\simeq N\xi\exp\left[\dfrac{\sum(\Delta\epsilon_{\rm hop}+f_{\rm osci})-f_{\rm trans}}{k_{\mathrm{B}}T}\right]\ ,$ (C1)
The summation encompasses all neighboring sites, while $f_{\rm osci}$ and $f_{\rm trans}$ represent the free energy contributions from oscillatory and translational motions, respectively, as defined in equation (A3). For carbonaceous substrates, the hopping energy barrier $\Delta\epsilon_{\rm hop}>0$ induced by each neighbor is quantified through the energy difference between adsorption configurations with zero and one neighbors, as tabulated in Table 1. In contrast, silicate substrates exhibited significantly weaker inter-adsorbate interactions relative to their substantial chemisorption energies, permitting the approximation of the hopping barrier simply by the absolute adsorption energy $|\Delta\epsilon_{\mathrm{ad}}|$.
Desorption. Molecular detachment from surface sites is characterized by a desorption rate governed by the expression:
$k_{\rm des}\simeq\nu\exp\left(\dfrac{\Delta\epsilon_{\mathrm{ad}}+\delta\epsilon_{\perp}-f_{\rm trans}}{k_{\mathrm{B}}T}\right)\ ,$ (C2)
where the vibrational frequency $\nu$ of adsorbed molecules corresponds to the parameter defined in equation (A2). The desorption energy barrier $\Delta\epsilon_{\mathrm{ad}}$ is computed using adsorption energies from Tables 1 and 2, with appropriate adjustments accounting for both the quantity and chemical identity of neighboring adsorbates to accurately capture cooperative effects in molecular binding.